OpenStudy (josd324life):

I just got switched into a new geometry class and I don't understand anything. Please help me.

OpenStudy (josd324life):

I attached the problem

OpenStudy (campbell_st):

These 3 questions all involve Pythagoras' Theorem \[c^2 = a^2 + b^2\] so in the 1st question you have \[50^2 = (2x)^2 + (x)^2\] square both terms on the right and you get \[2500 = 4x^2 + x^2\] now you can collect like terms and solve for x. the other questions work in the same way. In all 3 questions, the given measurement is on hypotenuse. hope it helps